OpenStudy (anonymous):

Find the exact value of the following: [\sin(-45)\]

OpenStudy (anonymous):

\[\sin(-45)\]

OpenStudy (anonymous):



OpenStudy (anonymous):

sin is opposite over hypotenuse

OpenStudy (anonymous):



OpenStudy (anonymous):

so \[\sin (-45)\]is \[\sqrt{2}\]

OpenStudy (anonymous):

\[\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

but its down so its negative

OpenStudy (anonymous):

\[-\sqrt{\frac{ 2 }{ 2 }}\]